ina multiple regression with six predictors in a sample of 67 cities what would be the critical value foran f test of overall significance at x .05
Answer #1

df regression = no. of precitors = 6

df total = N-1 = 67-1 = 66

df error = df total -df regression = 66-6 = 60

so, F critical value at α=0.05, df1 = 6 , df2 = 60 is  2.254 (answer)

{excel function for F critical value "=F.INV.RT(0.05,6,60) " }